Determine the mean, variance and standard deviation of the probability mass fucntion:

X: 0,1,2,3,4

P(x): 0.1, 0.2, 0.3, 0.3, 0.1


thanks!!!


1
Expert's answer
2022-03-03T08:24:07-0500

"M(X)=0*0.1+1*0.2+2*0.3+3*0.3+4*0.1=2.1"

"V(X)=M(X^2)-M^2(X)=0^2*0.1+1^2*0.2+2^2*0.3+3^2*0.3+4^2*0.1-2.1^2=5.7-4.41=1.29"

"\sigma(X)=\sqrt{V(X)}=\sqrt{1.29}\approx1.136"